With the assumptions of the Ricardian model, suppose the Home Country can produce 100 clothes, utilizing 25 workers when completely specialized, while the Foreign Country can produce 100 clothes, utilizing 20 workers when completely specialized. In this case, each country gains 1 additional unit of labor, which country gains more in the production of clothes from that additional unit and why? (H=Home, F=Foreign, C=Clothes)

The Foreign Country; because the MPL of clothes in foreign=5

The Home and Foreign Country gain the same number of units T

he Home Country; because the MPL of clothes at home=4

The Home Country; because the MPL of clothes at home=5
Answer #1

Marginal Product of labor in home=MPLH=100/25=4

Marginal Product of labor in foreign =MPLF=100/20=5

It is clear that MPLF>MPLH. So, Foreign will produce more clothes as compared to Home if one unit of extra labor is provided to each.

Correct option is

The Foreign Country; because the MPL of clothes in foreign=5